A bird is flying with velocity 5i+6j wrt wind and wind is blowing along x-axis with velocity u.If bird starts moving from A and after sometime reaches point B with AB making an angle of 37 deg with x axis.Find the value of u.

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Aparna, 

Net velocity= 5i+6j+ui=(5+u)i+6j, (5+u)/6=cot37,u=6cot37-5=2.96

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras